OpenStudy (anonymous):

Write 7/8 as an equivalent fraction with 56 as its denominator.

OpenStudy (anonymous):

What's 56 / 8?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

You can't divide 56 / 8? Put it in your calculator.

OpenStudy (anonymous):

its 7

OpenStudy (anonymous):

Yes, so we multiply the numerator and denominator by 7.

OpenStudy (anonymous):

\(\dfrac{7}{8} \rightarrow \dfrac{7 \times 7}{8 \times 7}\) So what's 7 * 7?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

Yes, that gives us \(\dfrac{49}{56}\).

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@inowalst What..?

OpenStudy (inowalst):

Um.. yeah. That was meant for the top, but lag.. >.> Sorry.